OpenStudy (anonymous):

What is the probabiltiy that you will spin a spinner numbered 1-6 and get a 4 three times in a row??

OpenStudy (anonymous):

1/6 for one time 1/12 for two 1/18 for three

OpenStudy (anonymous):

please medal

OpenStudy (anonymous):

I am confused Help

OpenStudy (campbell_st):

P(4) = 1/6 so the probability of a 4 three times P(4, 4, 4) = 1/6 x 1/6 x 1/6 =1/216

OpenStudy (anonymous):

That is so not right

OpenStudy (anonymous):

ok i will explain better out of 6 numbers one is a 4so one spin to land on 4 would be a 1/6 (one out of six chance) then you get two spins which means the 6 numbers get doubled so now you have a one out of 12 chance to hit four again but you want three spins in a row to land on four so you add another six which now makes your chances of spinning a 4 three times in a row one in 18 or 1/18

OpenStudy (campbell_st):

actually it is right... each spin is independent of the next. draw a tree diagram to represent each spin